OpenStudy (anonymous):

You have a job that pays $ 6.20 an hour. You get $ 9.50 an hour for overtime (anything over 40 hours). How many hours do you need to work to earn $452.25?

OpenStudy (callisto):

maximum salary you can earn without working overtime = 6.20*40 = $248.0 Let x be the number of hours working overtime 6.20*40 + 9.50x = 452.25 9.50x = 452.25 -6.20*40 9.50x = 204.25 x = 204.25 / 9.50 = 21.5 hours So total number of hours required = 21.5 +40 = work it our yourself